### v1

# Ab initio Exact Calculation of Strongly-Correlated Nucleonic Matter

Abstract
Dense nucleonic matter is of vital importance for understanding compact stars and inferring the transition into deconfined quark phase. We present the ab initio exact calculations of infinite nucleonic matter with the state-of-the-art full configuration-interaction quantum Monte Carlo (FCIQMC) method, enabling us to rigorously benchmark many-body methods and assess the degree to which the nucleonic matter is strongly correlated. This method has been numerically testified to exact diagonalization within a small basis. Calculations of nucleonic matter using the chiral nuclear forces reveal that the symmetric nuclear matter is strikingly strongly correlated than expected, raising questions on previous ab initio calculations of nuclear matter with truncations and offering insights into simultaneous descriptions of finite nuclei and infinite nucleonic matter from first principles.
Introduction.— Ab initio nuclear theory has made significant progress over the past two decades Epelbaum et al. (2020); Hergert (2020), however, exact many-body calculation of infinite nucleonic matter is still a missing piece. In the ab initio nuclear paradigm, starting from realistic nuclear forces, the full-configuration no-core shell model (NCSM) can be seen as exact diagonalization calculations of finite nuclei with active degrees of freedom of all nucleons Barrett et al. (2013). But NCSM is limited to light nuclei up to 16O owing to formidable computing costs. Moreover, it is essential to examine the applicability of ab initio calculations to heavy nuclei considering the in-medium effect Morris et al. (2018); Arthuis et al. (2020); Gysbers et al. (2019); Binder et al. (2014); Miyagi et al. (2022); Hu et al. (2024); Sammarruca and Millerson (2019), although realistic nuclear forces fitted to free-space scattering experiments are verified in few-body systems Entem and Machleidt (2003); Roth et al. (2005). To this end, ab initio calculations have to reproduce the empirical properties of infinite nuclear matter at the saturation density Simonis et al. (2017); Drischler et al. (2019); Jiang et al. (2020). However, most ab initio nuclear methods fail to describe finite nuclei and nuclear matter simultaneously Machleidt and Sammarruca (2024); Hebeler (2021). From this perspective, ab initio exact calculations of infinite nuclear matter are imperative.
The accurate predictions of equation of state (EoS) of cold dense nucleonic matter that are not accessible by terrestrial experiments is vital for understanding compact neutron stars in the multi-message astrophysical era Özel and Freire (2016); Reed et al. (2021); Essick et al. (2021); Neill et al. (2023); Annala et al. (2023). It also provides a stepping stone for inferring the phase transition scenario from nuclear matter to quark matter, informing us the possible existence of hybrid stars and non-perturbative properties of strong interactions Weih et al. (2020); Zhang and Li (2023); Fujimoto et al. (2023); Huang et al. (2022); Kedia et al. (2022). The theoretical challenge is related to the knowledge that nuclear matter are strongly correlated systems, in particular the symmetric nuclear matter (SNM), while the pure neutron matter (PNM) are relatively weakly correlated Gandolfi et al. (2015); Piarulli et al. (2020); Lovato et al. (2022); Drischler et al. (2020). This is indicated by previous studies that different ab initio methods have obvious discrepancies in SNM using hard nuclear forces Carbone et al. (2013); Rrapaj et al. (2016); Hagen et al. (2014); Baardsen et al. (2013); Marino et al. (2024); Zhen et al. (2025); Udiani (2025). Therefore, it will be interesting to see how the truncations applied in ab initio calculations of finite nuclei play in infinite nuclear matter.
It is impossible to perform exact diagonalization calculations of infinite nuclear matter by brute force Hjorth-Jensen et al. (2017a). There are many-body methods with truncations on high-order correlations applied to nuclear matter, including the many-body perturbation theory (MBPT) up to the 4th order Tichai et al. (2020); Drischler et al. (2019); Keller et al. (2023), coupled-cluster (CC) truncated at doublet amplitude level with perturbative triple corrections CCD(T) Hagen et al. (2014); Baardsen et al. (2013), self-consistent Green’s function (SCGF) truncated to the 3rd-order propagator Raimondi and Barbieri (2018); Barbieri et al. (2022); Marino et al. (2024), and the in-medium similarity renormalization group (IMSRG) truncated to the normal-ordered two-body level Udiani (2025); Zhen et al. (2025). On the other hand, various quantum monte carlo methods, such as auxiliary field diffusion Monte Carlo (AFDMC) Wlazłowski et al. (2014); Lynn et al. (2016); Piarulli et al. (2020) and configuration-interaction Monte Carlo (CIMC) Roggero et al. (2014); Arthuis et al. (2023), can in principle include high-order correlations but suffer the sign problem. To address the aforementioned difficulties, we hereby introduce the state-of-the-art full configuration-interaction quantum Monte Carlo (FCIQMC) method for ab initio exact calculations of infinite nuclear matter.
FCIQMC was originally developed in quantum chemistry and quickly has been widely used in molecular and condensed matter physics Booth et al. (2009); Shepherd et al. (2012); Booth et al. (2013); Cleland et al. (2010). FCIQMC has emerged as one of the most accurate many-body methods, particularly for strongly correlated systems Schwarz et al. (2015); Williams et al. (2020). Different from post-Hartree-Fock methods, FCIQMC is a projector QMC method that stochastically samples the ground-state wavefunction in the full Hilbert space. This approach also offers distinct advantages over other QMC techniques Spencer et al. (2019); Guther et al. (2020). Being formulated in a discrete configuration space, it is not restricted to local potentials and can naturally handle nonlocal forces from chiral effective field theory (EFT). Furthermore, FCIQMC employs a delicate walker annihilation algorithm to overcome the fermion sign problem in QMC methods Booth et al. (2009); Booth and Alavi (2010). This contrasts with methods like AFDMC or CIMC, which require the fixed-node approximation or guiding wavefunctions, introducing a bias that is difficult to assess or eliminate Zhang et al. (1997); Malone et al. (2020); Spencer et al. (2019); Guther et al. (2020). Therefore, as a method that in-principle is exact and possessing well-controlled systematic errors, FCIQMC holds the promise of providing benchmarking and advancing our understandings of dense nuclear matter and nuclear forces.
Methodology.— FCIQMC solves the imaginary-time Schrödinger equation Hammond et al. (1994),
where is the many-body Hamiltonian. In the long-time limit (), the wavefunction stochastically projects onto the exact ground state of .
The core idea of FCIQMC is to represent the wavefunction as a dynamic population of discrete, signed walkers distributed across the full configuration interaction (FCI) basis of all Slater determinants Booth et al. (2009), . The coefficient for each determinant is sampled by the number of walkers residing on it. The evolution of these walker populations is governed by the master equation:
where are the Hamiltonian matrix elements and is an energy offset known as the “shift”. This master equation is simulated via a stochastic algorithm applied at each time step , which comprises three steps Booth and Chan (2012):
1.
Spawning: Walkers on a determinant can “spawn” new walkers onto connected determinants (where ). The probability of creating a new walker is proportional to the magnitude of the off-diagonal matrix element .
2.
Death/Cloning: The diagonal term, , governs a local process where walkers on are removed (death) or duplicated (cloning), depending on whether the local energy is greater or less than the shift .
- 3.
The shift is dynamically adjusted to maintain a stable total walker number, . Once the simulation reaches equilibrium, i.e., fluctuates around a stable value, the ground-state energy is calculated using the projected estimator,
where is a reference determinant. The computational cost is linearly related to the total walker number. The ground-state energy and its statistical uncertainty are determined through reblocking analysis Flyvbjerg and Petersen (1989); Spencer .
To enhance the efficiency and stability for large systems, we employ the initiator approximation Cleland et al. (2010). This widely-used adaptation designates determinants with a walker population above a certain threshold as “initiators”. Spawning from non-initiator determinants is then restricted to only those determinants that are already occupied. This curtails the proliferation of noise from sparsely populated regions of the Hilbert space. While the initiator approximation introduces a systematic bias, such a bias is progressively reduced by increasing the total number of walkers, allowing for the extrapolation to exact solutions, the unbiased FCI limit Ghanem et al. (2021); Brand et al. (2022).
Results.—To validate our implementation and demonstrate the powerful capability of FCIQMC, we first benchmark it against several other advanced many-body methods using the exactly solvable Richardson pairing model Richardson (1963); Dukelsky et al. (2004). This model serves as an ideal testing ground for assessing the accuracy of many-body methods across a wide range of interaction strengths, from weak to strong couplings Rigo et al. (2023); Companys Franzke et al. (2024); Hjorth-Jensen et al. (2017b); Brolli et al. (2025). The Hamiltonian is given by
where is the single-particle level spacing (set to 1.0) and is the pairing strength. Our benchmark is performed for a system of nucleons in levels (a half-filled case), using a total walker population of for the FCIQMC calculations.
Figure 1 presents the correlation energies using different many-body methods and their deviations from the exact solution by direct diagonalzing. As shown, perturbative approaches like MBPT diverge rapidly as the interaction strength increases. While non-perturbative methods such as CCD, truncated SCGF using the algebraic diagrammatic construction [ADC(3)-D], and IMSRG(2) offer significant improvements in the weak-coupling regime, they still exhibit substantial deviations in the strong-coupling regime . In contrast, the FCIQMC results are in remarkable agreement with the exact energies across the entire range of , with statistical errors being almost negligible. The benchmarking shows the superior accuracy of FCIQMC, especially in strongly coupled situations where truncated many-body methods break down (see Fig. 1). Our implementation for the Richardson model is made publicly available for reproducibility and further development Hu .
We now turn to much more challenging calculations of the realistic system of infinite nuclear matter. The many-body Hamiltonian, , includes the kinetic energy term alongside interaction terms corresponding to nucleon-nucleon (NN) and three-nucleon forces (3NF) that are derived from EFT. To compare with other methods, the 3NF is included at the normal-ordered two-body level Hebeler and Schwenk (2010); Hagen et al. (2014), although FCIQMC can naturally deal with the full 3NF. The many-body basis is constructed by Slater determinants built with single-particle (SP) momentum eigenstates, on the lattice in momentum space. Our calculations are performed in a cubic box of volume with nucleons at a density , employing periodic boundary conditions which discretize the momenta with a spacing of . The SP basis is defined by a momentum cutoff , which should be sufficiently large to ensure convergence of solutions. Two distinct families of EFT forces at the next-to-next-to-leading-order (N2LO) are used: N2LO(Hüther) Hüther et al. (2020) and Ekström et al. (2018); Jiang et al. (2020), where the later explicitly includes the isobar degree of freedom.
Before proceeding to large-scale calculations with FCIQMC, we perform a rigorous benchmark for symmetric nuclear matter (SNM) in a small model space where exact results can be obtained via direct diagonalization, compared with MBPT and IMSRG(2) methods. The results are shown in Fig. 2. Across all densities and for both types of nuclear forces, FCIQMC calculations with walkers reproduce exact energies with extraordinary accuracy, with deviations being orders of magnitude smaller than other methods. In contrast, both MBPT and IMSRG(2) exhibit significant discrepancies that grow substantially with increasing density and particularly for the harder interactions. This benchmark not only validates our FCIQMC implementation for infinite nuclear matter but also provides a clear illustration of the limitations of these widely-used truncated schemes for infinite nucleonic matter.
Next, we present large-scale calculations with the -full interaction, (450 MeV cutoff) Jiang et al. (2020), which has been successful in describing binding energies and radii of finite nuclei up to 132Sn Koszorús et al. (2021); Sommer et al. (2022); Novario et al. (2023); Nies et al. (2023); König et al. (2024); Door et al. (2025) but whose performance for nuclear matter saturation has been debated Ekström et al. (2024); Marino et al. (2024).
As shown in Fig. 3, for PNM, a relatively weakly correlated system, all methods yield similar energies per particle with differences below 1 MeV. The estimated uncertainties from EFT truncation at N2LO are also shown Epelbaum et al. (2015); Binder et al. (2016, 2018). For the strongly-correlated SNM, however, the picture changes dramatically. Our FCIQMC calculations predict saturation properties that are in good agreement with the empirical saturation window fm-3 and MeV Drischler et al. (2016, 2019). The resulting symmetry energy is also consistent with experimental constraints Li et al. (2019). In contrast, both MBPT and IMSRG(2) calculations predict significant overbinding at higher densities.
Since FCIQMC incorporates correlations to all orders, the discrepancy between FCIQMC and other methods provides quantitative evidence for the crucial role of high-order many-body correlations. As illustrated in Fig. 3(b), the energy contribution from the missing correlations is about 4 MeV around saturation. This informs us strikingly that SNM is much more strongly-correlated than expected. The high-order contribution is of the same magnitude as the estimated EFT truncation uncertainty at N2LO. This implies that for accurate descriptions of nuclear matter, a precise many-body solver and the chiral expansion to higher-order are equally important. Our study confirms that the force can indeed provide a consistent description of both finite nuclei and infinite nuclear matter, but only when many-body correlations are exactly handled.
We also perform a parallel analysis with the -less N2LO interaction, which is able to reproduce experimental binding energies and charge radii of medium-mass nuclei up to 78Ni Hüther et al. (2020). Previous MBPT calculations concluded that this interaction fails to describe SNM saturation Sammarruca and Millerson (2021); Machleidt and Sammarruca (2024). Our FCIQMC calculations, shown in Fig. 4, find that this conclusion relies on many-body methods being used.
Our results show that perturbative methods like MBPT(3) severely overestimate the binding energy of SNM with the -less N2LO interaction compared to FCIQMC results. The non-perturbative IMSRG(2), while yielding a fortuitous saturation point near the empirical value, is still in disagreement with the FCIQMC benchmark. This highlights a critical issue that the many-body uncertainty arising from methodological truncations can be so large, especially for harder interactions, that it completely obscures the true properties of the underlying Hamiltonian. Here the discrepancy between FCIQMC and other methods is significantly larger than that using the -full force, demonstrating that the high-order correlations become increasingly important for hard nuclear forces. This finding underscores an important message to the field: to reliably assess the quality of a chiral interaction and to make credible progress in solving the long-standing saturation problem, it is necessary to improve the truncation schemes of existing post-Hartree-Fock methods. In this respect, FCIQMC is useful to disentangle the potential deficiencies of a nuclear Hamiltonian from the many-body approximations.
Summary.— In this Letter, we introduce the state-of-the-art FCIQMC method for ab initio studies of infinite nuclear matter. By stochastically sampling in the complete many-body Hilbert space, FCIQMC can include full many-body correlations, circumventing both the truncation errors of methods like CC and IMSRG. FCIQMC also provides a solution to the notorious sign problem in QMC methods and avoids the fixed-node biases of other QMC approaches. This work establishes FCIQMC as a rigorous benchmarking tool for infinite nuclear matter, which is one of the most challenging strongly correlated systems in nature.
Our primary finding is that high-order many-body correlations, largely neglected in other methods, play a crucial role in describing nuclear matter saturation. For modern N2LO chiral interactions, the energy contribution from these correlations is found to be several MeV, a magnitude comparable to the theoretical uncertainty of the chiral expansion itself, indicating that SNM is much more strongly correlated than expected. The present work also demonstrates a significant advance in resolving the long-standing problem to simultaneously describe the bulk properties of finite nuclei and infinite nuclear matter from a single Hamiltonian with the -full interaction. The FCIQMC as a robust tool to predict nuclear EoS establishes a reliable link between microscopic nuclear forces and the astrophysical observables of neutron stars.
The capability of FCIQMC to controllably capture correlations to all orders opens new frontiers in ab initio nuclear physics. It enables a clear separation of Hamiltonian versus many-body uncertainties, which is crucial for guiding the development of the next generation of high-precision interactions. Furthermore, applying FCIQMC to other frameworks will greatly advance ab initio calculations of finite nuclei, nuclear reactions, and electroweak processes, with considerable impacts on our understanding of phenomena across nuclear physics and astrophysics.
The authors would like to thank J. G. Li and S. Zhang for useful discussions. This work is supported by the National Key R&D Program of China under Grant No. 2023YFA1606401 and 2024YFA1610900; the National Natural Science Foundation of China under Grants No. 12335007, 12035001, 12475118 and 11921006. Numerical computations were performed at the High-Performance Computing Platform of Peking University.