### v3

# Ab initio Exact Calculation of Strongly-Correlated Nucleonic Matter

Abstract
Dense nucleonic matter is of vital importance for understanding compact stars and inferring the transition into deconfined quark phase. We present ab initio exact calculations of infinite nucleonic matter with the state-of-the-art full configuration-interaction quantum Monte Carlo (FCIQMC) method, enabling us to rigorously benchmark many-body methods and assess the degree to which the nucleonic matter is correlated. This method has been numerically validated by exact diagonalization within a small model space. Calculations of nucleonic matter using chiral nuclear forces reveal that symmetric nuclear matter is strikingly strongly correlated, raising questions on previous ab initio calculations of nuclear matter with many-body expansion truncations and offering insights into simultaneous descriptions of finite nuclei and infinite nucleonic matter from first principles.
Introduction.— Ab initio nuclear theory has made significant progress over the past two decades Epelbaum et al. (2020); Hergert (2020), however, exact many-body calculation of infinite nucleonic matter is still a missing piece. In the ab initio nuclear paradigm, starting from realistic nuclear forces, the full-configuration no-core shell model (NCSM) can be seen as exact diagonalization calculations of finite nuclei with active degrees of freedom of all nucleons Barrett et al. (2013). But NCSM is limited to light nuclei up to 16O owing to formidable computing costs. Moreover, it is essential to examine the applicability of ab initio calculations to heavy nuclei considering the in-medium effect Morris and others (2018); Arthuis et al. (2020); Gysbers and others (2019); Binder et al. (2014); Miyagi et al. (2022); Hu and others (2024); Sammarruca and Millerson (2019), although realistic nuclear forces fitted to free-space scattering experiments are verified in few-body systems Entem and Machleidt (2003); Roth et al. (2005). To this end, ab initio calculations have to reproduce the empirical properties of infinite nuclear matter at the saturation density Simonis et al. (2017); Drischler et al. (2019); Jiang et al. (2020). However, most ab initio nuclear methods fail to describe finite nuclei and nuclear matter simultaneously Machleidt and Sammarruca (2024); Hebeler (2021). From this perspective, ab initio exact calculations of infinite nuclear matter are imperative.
The accurate prediction of equation of state (EoS) of cold dense nucleonic matter that is not accessible by terrestrial experiments is vital for understanding compact neutron stars in the multi-messenger astrophysical era Özel and Freire (2016); Reed et al. (2021); Essick et al. (2021); Neill et al. (2023); Annala et al. (2023). It also provides a stepping stone for inferring the phase transition scenario from nuclear matter to quark matter, informing us of the possible existence of hybrid stars and non-perturbative properties of strong interactions Weih et al. (2020); Zhang and Li (2023); Fujimoto et al. (2023); Huang and others (2022); Kedia et al. (2022). The theoretical challenge is related to the knowledge that nuclear matter is a strongly correlated system, in particular the symmetric nuclear matter (SNM), while the pure neutron matter (PNM) is relatively weakly correlated Gandolfi et al. (2015); Piarulli et al. (2020); Lovato et al. (2022); Drischler et al. (2020). This is indicated by previous studies that different ab initio methods have obvious discrepancies in SNM using hard nuclear forces Carbone et al. (2013); Rrapaj et al. (2016); Hagen and others (2014); Baardsen et al. (2013); Marino et al. (2024a); Zhen et al. (2025); Udiani (2025). Therefore, it will be interesting to see how the truncations applied in ab initio calculations of finite nuclei play in infinite nuclear matter.
It is impossible to perform exact diagonalization calculations of infinite nuclear matter by brute force Hjorth-Jensen et al. (2017a). There are ab initio methods with many-body expansion truncations applied to nuclear matter, including the many-body perturbation theory (MBPT) up to the 4th order Tichai et al. (2020); Drischler et al. (2019); Keller et al. (2023), coupled-cluster (CC) truncated at the doubles excitation level with perturbative triple corrections CCD(T) Hagen and others (2014); Baardsen et al. (2013), self-consistent Green’s function (SCGF) theory Rios (2020) based on the algebraic diagrammatic construction (ADC) approximation scheme Raimondi and Barbieri (2018); Barbieri et al. (2022); Marino et al. (2024a, 2026), and the in-medium similarity renormalization group (IMSRG) truncated to the normal-ordered two-body level Udiani (2025); Zhen et al. (2025). While these basis-expansion techniques are highly effective with soft potentials, Quantum Monte Carlo (QMC) approaches are generally advantageous for handling harder interactions. In coordinate space, the auxiliary field diffusion Monte Carlo (AFDMC) Wlazłowski et al. (2014); Lynn et al. (2016); Piarulli et al. (2020) and recent variational calculations using neural-network quantum states Carleo and Troyer (2017); Pfau et al. (2020); Hermann et al. (2020); Adams et al. (2021); Keeble and Rios (2020); Luo and Clark (2019) have been developed. In configuration space, methods such as the configuration-interaction Monte Carlo (CIMC) Roggero et al. (2014); Arthuis et al. (2023) have also been available. Although these QMC methods can in principle include high-order correlations, they often suffer from the fermion sign problem or rely on the quality of variational ansatz. To address the aforementioned difficulties, we hereby introduce the state-of-the-art full configuration-interaction quantum Monte Carlo (FCIQMC) method for ab initio exact calculations of infinite nuclear matter.
FCIQMC was originally developed in quantum chemistry and quickly has been widely used in molecular and condensed matter physics Booth et al. (2009); Shepherd et al. (2012); Booth et al. (2013); Cleland et al. (2010). FCIQMC has emerged as one of the most accurate many-body methods, particularly for strongly correlated systems Schwarz et al. (2015); Williams and others (2020). Different from post-Hartree-Fock methods, FCIQMC is a projector QMC method that stochastically samples the ground-state wavefunction in the full Hilbert space. This approach also offers distinct advantages over other QMC techniques Spencer and others (2019); Guther and others (2020). Being formulated in a discrete configuration space, unlike coordinate-space methods such as AFDMC, FCIQMC is not restricted to local potentials and can naturally handle nonlocal forces from chiral effective field theory (EFT). Furthermore, FCIQMC employs a delicate walker annihilation algorithm to overcome the fermion sign problem in QMC methods Booth et al. (2009); Booth and Alavi (2010). This contrasts with methods like AFDMC or CIMC, which require the fixed-node approximation or guiding wavefunctions, introducing a bias that is difficult to assess or eliminate Zhang et al. (1997); Malone et al. (2020); Spencer and others (2019); Guther and others (2020). Therefore, as a method that in-principle is exact and possessing well-controlled systematic errors, FCIQMC holds the promise of providing benchmarking and advancing our understandings of dense nuclear matter and nuclear forces.
Methodology.— FCIQMC solves the imaginary-time Schrödinger equation Hammond et al. (1994),
where is the many-body Hamiltonian. In the long-time limit (), the wavefunction stochastically projects onto the exact ground state of .
The core idea of FCIQMC is to represent the wavefunction as a dynamic population of discrete, signed walkers distributed across the full configuration interaction (FCI) basis of all Slater determinants Booth et al. (2009), . The coefficient for each determinant is sampled by the number of walkers residing on it. The evolution of these walker populations is governed by the master equation:
where are the Hamiltonian matrix elements, and is an energy offset known as shift which is introduced to prevent the true ground state from decaying to zero. This master equation is simulated via a stochastic algorithm applied at each time step , which comprises three steps Booth and Chan (2012):
1.
Spawning: Walkers on a determinant can “spawn” new walkers onto connected determinants (where ). The probability of creating a new walker is proportional to the magnitude of the off-diagonal matrix element .
2.
Death/Cloning: The diagonal term, , governs a local process where walkers on are removed (death) or duplicated (cloning), depending on whether the local energy is greater or less than the shift .
- 3.
The shift is dynamically adjusted to maintain a stable total walker number, . Once the simulation reaches equilibrium, i.e., fluctuates around a stable value, the ground-state energy is calculated using the projected estimator,
where is a reference determinant. The computational cost is linearly related to the total walker number. The ground-state energy and its statistical uncertainty are determined through reblocking analysis Flyvbjerg and Petersen (1989); Spencer .
To enhance the efficiency and stability for large systems, we employ the initiator approximation Cleland et al. (2010). This widely-used adaptation designates determinants with a walker population above a certain threshold as “initiators”. Spawning from non-initiator determinants is then restricted to only those determinants that are already occupied. This curtails the proliferation of noise from sparsely populated regions of the Hilbert space. While the initiator approximation introduces a systematic bias, such a bias is progressively reduced by increasing the total number of walkers, allowing for the extrapolation to exact solutions, the unbiased FCI limit Ghanem et al. (2021); Brand et al. (2022). The adaptive-shift method Ghanem et al. (2019); Guther and others (2020) has been used to accelerate the convergence within large model spaces. The full details of the FCIQMC algorithms are shown in the Supplemental Material 1.
Results.—To validate our implementation and demonstrate the powerful capability of FCIQMC, we first benchmark it against several other advanced many-body methods using the exactly solvable Richardson pairing model Richardson (1963); Dukelsky et al. (2004). This model serves as an ideal testing ground for assessing the accuracy of many-body methods across a wide range of interaction strengths, from weak to strong couplings Rigo et al. (2023); Companys Franzke et al. (2024); Hjorth-Jensen et al. (2017b); Brolli et al. (2025). The Hamiltonian is given by
where is the single-particle level spacing (set to 1.0) and is the pairing strength. Our benchmark is performed for a system of nucleons in levels (a half-filled case), using a total walker population of for the FCIQMC calculations.
Figure 1 presents the correlation energies using different many-body methods and their deviations from the exact solution by direct diagonalzing. As shown, perturbative approaches like MBPT at order (=2, 3, 4) depart from exact solutions rapidly as the interaction strength increases. While non-perturbative methods such as CCD, ADC-SCGF at third-order in the ADC hierarchy plus CCD amplitude corrections [ADC(3)-D] Barbieri and Carbone (2017); Marino et al. (2026), and IMSRG(2) offer significant improvements in the weak-coupling regime, they still exhibit substantial deviations in the strong-coupling regime . In contrast, the FCIQMC results are in remarkable agreement with the exact energies across the entire range of , with statistical errors being almost negligible. The benchmarking shows the superior accuracy of FCIQMC, especially in strongly coupled situations where truncated many-body methods break down (see Fig. 1). Our implementation for the Richardson model is made publicly available for reproducibility and further development Hu .
We now turn to much more challenging calculations of the realistic system of infinite nuclear matter. The many-body Hamiltonian, , includes the kinetic energy term alongside interaction terms corresponding to nucleon-nucleon (NN) and three-nucleon forces (3NF) that are derived from EFT. To compare with other methods, the 3NF is included at the normal-ordered two-body level Hebeler and Schwenk (2010); Hagen and others (2014), although FCIQMC can naturally deal with the full 3NF. The many-body basis is constructed by Slater determinants built from single-particle (SP) momentum eigenstates, on the lattice in momentum space. Our calculations are performed in a cubic box of volume with nucleons at a density , employing periodic boundary conditions which discretize the momenta with a spacing of . The SP basis is defined by a momentum cutoff , which should be sufficiently large to ensure convergence of solutions. Two distinct families of EFT forces at the next-to-next-to-leading-order (N2LO) are used: N2LO(Hüther) Hüther et al. (2020) and Ekström et al. (2018); Jiang et al. (2020), where the latter explicitly includes the isobar degree of freedom.
Before proceeding to large-scale calculations with FCIQMC, we perform a rigorous benchmark for SNM in a small model space where exact results can be obtained via direct diagonalization, compared with MBPT and IMSRG(2) methods. The results are shown in Fig. 2. Across all densities and for both families of nuclear forces, FCIQMC calculations with walkers reproduce exact energies with extraordinary accuracy, with deviations being orders of magnitude smaller than other methods. In contrast, both MBPT and IMSRG(2) exhibit significant discrepancies that grow substantially with increasing density and particularly for the harder interactions. MBPT(3) yields reasonable agreement in some cases, which depends on the interaction being used. This benchmark not only validates our FCIQMC implementation for infinite nuclear matter but also provides a clear illustration of the limitations of these widely-used truncated schemes for infinite nucleonic matter.
Next, we present large-scale calculations with the -full interaction, (450 MeV cutoff) Jiang et al. (2020), which has been successful in describing binding energies and radii of finite nuclei up to 132Sn Koszorús and others (2021); Sommer and others (2022); Novario et al. (2023); Nies and others (2023); König and others (2024); Door and others (2025) but whose performance for nuclear matter saturation has been debated Marino et al. (2026); Ekström and others (2024); Marino et al. (2024b). We compare FCIQMC results with those from other five methods, namely MBPT(2), MBPT(3), CCD, ADC(3)-D and IMSRG(2). All the calculations were performed in the same model space which contains 66 (76) nucleons and 682 (1364) SP states for PNM (SNM). The number of SP states is sufficiently large for the basis convergence. The uncertainty analysis can be found in the Supplemental Material 1.
As shown in Fig. 3, for PNM, a relatively weakly correlated system, all methods yield similar energies per particle with differences below 0.5 MeV. For the SNM, however, the differences between FCIQMC and other methods become larger up to 2 MeV. Since SNM is more strongly correlated than PNM, the effect of the missing higher-order many-correlations is much more evident. All methods overestimate both the saturation density and the saturation energy compared to the empirical saturation window fm-3 and MeV Drischler et al. (2016, 2019). Nevertheless, the resulting symmetry energies are consistent with experimental constraints Li et al. (2019). FCIQMC gives the lowest binding energies among all methods. We find that results of MBPT(3) and ADC(3)-D are relatively close to that of FCIQMC.
We also perform a parallel analysis with the -less N2LO interaction Hüther et al. (2020) with the same cutoff of 450 MeV, which is able to reproduce experimental binding energies and charge radii of medium-mass nuclei up to 78Ni using IMSRG(2) Hüther et al. (2020). Previous Brueckner-Hartree-Fock calculations concluded that this interaction fails to describe SNM saturation Sammarruca and Millerson (2021b); Machleidt and Sammarruca (2024); Sammarruca and Millerson (2021a). Our calculations, shown in Fig. 4, find that this conclusion relies on many-body methods being used.
Our results show much more significant differences in the binding energies of SNM among different methods with the -less N2LO interaction. This is because this interaction is harder than the -full interaction in Fig. 3. The differences in PNM are again very small. For SNM, we see the MBPT(3) is close to FCIQMC, while the difference between MBPT(2) and MBPT(3) is so large that the convergence is not guaranteed. IMSRG(2) accidently reproduces the saturation point with this interaction. FCIQMC gives the lowest binding energies for SNM and the differences compared to other methods increase remarkably towards high densities.
Since FCIQMC incorporates correlations to all orders, the discrepancy between FCIQMC and other methods provides quantitative evidence for the crucial role of high-order many-body correlations. As illustrated in Fig. 4(b), the energy contribution from the missing correlations is about 10 MeV around saturation and about 40 MeV at 2.0. This informs us that SNM is much more strongly-correlated than expected at higher densities. The results also demonstrated that the high-order correlations become strikingly important for hard nuclear forces. The high-order contribution is so large that it could be more significant than the EFT truncation uncertainty Epelbaum et al. (2015); Binder and others (2016, 2018) at high densities. This finding underscores an important message to the field: to reliably assess the quality of a chiral interaction and to make credible progress in solving the long-standing saturation problem Akmal et al. (1998); Lejeune et al. (2000); Hebeler et al. (2011); Machleidt and Sammarruca (2024); Hebeler (2021), it is necessary to improve the truncation schemes of existing post-Hartree-Fock methods. In this respect, FCIQMC is useful to disentangle the potential deficiencies of a nuclear Hamiltonian from the many-body approximations.
Summary.— In this Letter, we introduce the state-of-the-art FCIQMC method for ab initio studies of infinite nuclear matter. By stochastically sampling in the complete many-body Hilbert space, FCIQMC can include full many-body correlations, circumventing the truncation errors of expansion methods like CC, IMSRG and ADC-SCGF. FCIQMC also provides a solution to the notorious sign problem in QMC methods and avoids the fixed-node biases of other QMC approaches. The wave functions are directly sampled in FCIQMC, facilitating the calculations of various observables. This work establishes FCIQMC as a rigorous benchmarking tool for infinite nuclear matter, which is one of the most challenging strongly correlated systems in nature.
Our primary finding is that high-order many-body correlations, largely truncated in other methods, play a crucial role in describing nuclear matter saturation. For the hard chiral interaction such as N2LO(Hüther) Hüther et al. (2020), the energy contribution from these correlations is found to be strikingly large up to 40 MeV at 2.0, a magnitude comparable to the theoretical uncertainty of the chiral expansion itself, indicating that SNM is much more strongly correlated than expected. By eliminating many-body uncertainties, this work makes a substantial step in resolving the long-standing problem to simultaneously describe the bulk properties of finite nuclei and infinite nuclear matter from a single Hamiltonian. The FCIQMC as a robust tool to predict nuclear EoS also establishes a reliable link between microscopic nuclear forces and the astrophysical observables of neutron stars.
The capability of FCIQMC to controllably capture correlations to all orders opens new frontiers in ab initio nuclear physics. It enables a clear separation of Hamiltonian versus many-body uncertainties, which is crucial for guiding the development of the next generation of high-precision interactions. Furthermore, applying FCIQMC to other frameworks will greatly advance ab initio calculations of finite nuclei, nuclear reactions, and electroweak processes, with considerable impacts on our understanding of phenomena across nuclear physics and astrophysics.
Acknowledgments.— We are grateful to A. Alavi for valuable suggestions and discussions on the FCIQMC algorithm, and thank J. G. Li, S. Zhang, Carlo Barbieri and Gianluca Colò for useful discussions. This work has been supported by the National Key R&D Program of China under Grant No. 2023YFA1606401 and 2024YFA1610900; the National Natural Science Foundation of China under Grants No. 12335007, 12035001, 12475118 and 11921006, and the High-Performance Computing Platform of Peking University. F.M was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 279384907 – SFB 1245, and through the Cluster of Excellence “Precision Physics, Fundamental Interactions, and Structure of Matter” (PRISMA+ EXC 2118/1, Project ID 39083149). We acknowledge the CINECA awards AbINEF (HP10B3BG09) and RespGF (HP10BQMECT) under the ISCRA initiative, for the availability of highperformance computing resources and support. This work used the DiRAC Data Intensive service (DIaL3) at the University of Leicester, managed by the University of Leicester Research Computing Service on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The DiRAC service at Leicester was funded by BEIS, UKRI and STFC capital funding and STFC operations grants. DiRAC is part of the UKRI Digital Research Infrastructure.
Data availability.—The data that support the findings of this article and source code of FCIQMC for the pairing model are openly available Hu .